OpenStudy (henrietepurina):

Algebra 1 question... Medal + Fan :)

OpenStudy (henrietepurina):

Write the expression x2 + 5x + 6 as a product of two linear expressions. Show your work and justify each step.

OpenStudy (henrietepurina):

@asnaseer @confluxepic @Compassionate

OpenStudy (asnaseer):

If you want to learn about how to factor quadratic expressions then I can recommend these two sites: 1) https://www.khanacademy.org/math/algebra/multiplying-factoring-expression/factoring-quadratic-expressions/v/factoring-quadratic-expressions 2) http://www.mathsisfun.com/algebra/factoring-quadratics.html

OpenStudy (henrietepurina):

So am I suppose to factor then to get the answer?

OpenStudy (henrietepurina):

@asnaseer ?

OpenStudy (asnaseer):

The question asks "...as a product of two linear expressions" The expression you are given: \(x^2 + 5x + 6\) is not a linear expression - it is a quadratic expression. This quadratic expression can be factorised as a product of two linear expressions. Follow the links I gave you to see how this is achieved.

OpenStudy (michele_laino):

please , note that your function can be factorized as follows: \[x ^{2}+5x+6=(x-x _{1})(x-x _{2})\] where x_1 and x_2 are the solution of this equation: \[x ^{2}+5x+6=0\]